# Probabilistic approach to feedback control enhances multi-legged locomotion on rugged landscapes

### v1

@@ V-B Vertical motion reduces environmental disturbance
Figure 10: Empirical evidence supporting the necessity of vertical motion modulation. A. Empirical results demonstrate the relationship between the robot’s speed $\bar{v}$ and the vertical amplitude $A_{v}$. Here, $N$ represents the number of leg pairs on the robot. For example, if $N=4$, the robot has 8 legs. B. Snapshots show the differences in speed for a 12-legged robot with varying vertical amplitudes $A_{v}$.
@@
To verify our hypothesis through more real-world experiments, we tested our 12-legged robot on terrain with varying rugosity using different vertical amplitudes $A_{v}$. A motion tracking system and binary contact sensors were used to monitor the robot’s speed and contact ratio $\gamma$, respectively. The results (Fig. 9.A) show as $A_{v}$ increases, both the robot’s speed and contact ratio $\gamma$ tend to concentrate around the central area near (0.5, 0.5). These findings suggest that vertical body undulation effectively helps the robot mitigate environmental disturbances, improving both its speed and contact state.
@@ VI Enhance Robot Performance on Rugged Landscapes Through Feedback Control
In Section V, we show how vertical body motion effectively mitigates environmental disturbances on the robot’s speed. Our experimental findings, depicted in Fig. 10, highlight the importance of adapting vertical motion to optimize speed in response to changes in terrain rugosity. Specifically, the empirical data suggests that the robot attains maximum forward speed by adjusting vertical amplitudes ($A_{v}$) according to the terrain’s rugosity.
@@ VI-A Vertical motion adaptation based feedback controller
Figure 12: Controller comparison in laboratory experiments. Experiments were conducted on 12-legged (N = 6) and 8-legged (N = 4) robots to compare the performance of open-loop and feedback controllers. Here, D represents the forward displacement of the robots.
@@ VI-B Lab based experiment
The feedback controller’s efficacy was assessed through experimentation conducted on laboratory-based rough terrain, specifically focusing on terrain with higher rugosity ($R_{g}=0.32$). Over seven motion cycles, the contact state of each leg and robot’s forward displacement were recorded. Fig. 11.A.2 shows the temporal evolution of the measured $\gamma_{s}$ alongside the corresponding $A_{v}$ for a representative trial of the feedback control experiments.
@@
In previous feedback experiments, the robot adjusted its vertical amplitude on a per-cycle basis. To determine if its modulation frequency was optimal, we conducted additional tests on both 8-legged and 12-legged configurations with varying modulation frequencies. Our empirical findings suggest that cycle-wise vertical motion modulation emerges as the optimal approach. This method not only enhances the robot’s performance on rough terrain but also ensures computational efficiency.
@@ VI-C Outdoor experiment
We tested the feedback controller’s performance on five outdoor terrains with varied obstacles, including tree debris, grass, boulders, mud and rocks (Fig.14). In these highly unpredictable environments, the feedback controller increased the robot’s speed by 30-60%, reaching up to 60% of its maximum velocity in open areas. Additional details and visual demonstrations of the outdoor experiments can be accessed in the Supplementary Information (SI) videos.
@@ VII Conclusion
However, our analysis contained errors due to two key assumptions, along with other factors. First, we assumed that when a leg loses contact, the support force is lost rather than redistributed to other legs, based on the hypothesis that the force transfers to the robot’s belly. Second, we assumed the probability distribution of the slipping angle on flat terrain remains constant, though it may change on rough terrain. Additionally, sensor noise introduces errors in measuring the contact ratio, $\gamma$. To simplify calculations, we neglected the impact of collisions between the robot and the environment, despite their effect on thrust generation. These factors contributed to approximately 15% of the data points falling outside our predictions (Fig.9.B).
@@
We employed a simple sinusoidal traveling wave model to characterize the robot’s vertical body undulation. While effective in mitigating environmental disturbances through amplitude adjustments, this approach cannot generate more complex vertical body shapes. Future research will investigate the impact of more discrete vertical body shape on robot performance, focusing on how subtle variations can optimize locomotion efficiency and adaptability on complex terrains.
@@
Appendix A Motion track for experiments
@@ Acknowledgment
We would like to thank Daniel Irvine and Jeff Aguilar for the helpful discussion.

### v2

@@ V-B Vertical motion reduces environmental disturbance
To verify our hypothesis through more real-world experiments, we tested our 12-legged robot on terrain with varying rugosity using different vertical amplitudes $A_{v}$. A motion tracking system and binary contact sensors were used to monitor the robot’s speed and contact ratio $\gamma$, respectively. The results (Fig. 9.A) show as $A_{v}$ increases, both the robot’s speed and contact ratio $\gamma$ tend to concentrate around the central area near (0.5, 0.5). Specifically, on terrain with $R_{g}=0.17$, the optimal vertical amplitude $A_{v}$ increases the average contact ratio $\gamma$ by 25% and speed by 30%, compared to the gait without vertical wave ($A_{v}=0$). On terrain with $R_{g}=0.32$, the optimal $A_{v}$ increases the average $\gamma$ by 18% and speed by $20\%$. These findings suggest that vertical body undulation effectively helps the robot mitigate environmental disturbances, improving both its speed and contact state.
@@ VI Enhance Robot Performance on Rugged Landscapes Through Feedback Control
Figure 10: Empirical evidence supporting the necessity of vertical motion modulation. A. Empirical results demonstrate the relationship between the robot’s speed $\bar{v}$ and the vertical amplitude $A_{v}$. Here, $N$ represents the number of leg pairs on the robot. For example, if $N=4$, the robot has 8 legs. We conducted 10 experimental trials for each data point. Error bars represent the standard deviation. B. Snapshots show the differences in speed for a 12-legged robot with varying vertical amplitudes $A_{v}$.
In Section V, we show how vertical body motion effectively mitigates environmental disturbances on the robot’s speed. Our experimental findings, depicted in Fig. 10, highlight the importance of adapting vertical motion to optimize speed in response to changes in terrain rugosity. Specifically, the empirical data suggests that robot’s speed $\bar{v}$ is reduced by increasing vertical amplitude $A_{v}$ on flat ground ($R_{g}=0$). However, on rough terrain ($R_{g}=0.17\text{ or }0.32$), increasing $A_{v}$ can increase $\bar{v}$. These findings indicate that implementing feedback control of $A_{v}$ based on terrain rugosity could be beneficial for maximizing the robot’s forward speed. The sensitivity of the robot’s performance to temporal frequency is discussed in the Appendix. -E.
@@ VI-A Vertical motion adaptation based feedback controller
Figure 12: Controller comparison in laboratory experiments. Experiments were conducted on 12-legged (N = 6) and 8-legged (N = 4) robots to compare the performance of open-loop and feedback controllers. Here, D represents the forward displacement of the robots. Each condition was tested in 10 trials, and the shaded region represents the standard deviation.
Figure 13: Controller comparison in composite terrain. The left plot shows the forward displacement over time for both open-loop and feedback controllers as the robot traverses a composite terrain consisting of two roughness levels ($R_{g}=0.17$ and $R_{g}=0.32$). Shaded regions represent the standard deviation over five trials. The right image shows the 12-legged robot navigating the composite terrain, which features abrupt transitions in elevation and surface characteristics.
@@ VI-B Lab based experiment
The feedback controller’s efficacy was assessed through experimentation conducted on laboratory-based rough terrain, specifically focusing on terrain with higher rugosity ($R_{g}=0.32$). Over seven motion cycles, the contact state of each leg and robot’s forward displacement were recorded. Fig. 11.A shows the temporal evolution of the measured $\gamma_{s}$ alongside the corresponding $A_{v}$ for a representative trial of the feedback control experiments. We tuned $K_{p}$ from 0 to 180 degrees to obtain the optimal performance for the feedback controller. We set $K_{p}$ as 90 degrees for the lab-based experiments. We also conducted experiments for $K_{p}$ sensitivity analysis and the results can be found in Appendix. -F.
@@
To evaluate the real-time performance of the feedback controller, we tested the robot on a composite terrain combining sections of $R_{g}=0.17$ and $R_{g}=0.32$ (Fig. 13). The rapid transition in terrain rugosity at the junction of these two surfaces presents additional challenges for the controller. We ran the robot for eight motion cycles on this composite terrain using both open-loop and feedback controllers, conducting five trials for each condition. As shown in Fig. 13, the feedback controller improved the robot’s speed by approximately 50%, demonstrating its effectiveness even under rapidly changing terrain conditions.
In previous feedback experiments, the robot adjusted its vertical amplitude on a per-cycle basis. To determine if its modulation frequency was optimal, we conducted additional tests on both 8-legged and 12-legged configurations with varying modulation frequencies. Our empirical findings (Fig.14) suggest that cycle-wise vertical motion modulation emerges as the optimal approach. This method not only enhances the robot’s performance on rough terrain but also ensures computational efficiency.
@@ VI-C Outdoor experiment
We tested the feedback controller’s performance on five outdoor terrains with varied obstacles, including tree debris, grass, boulders, mud and rocks (Fig.15). Terrain a consists of a mixture of robot-sized boulders, leaves, and mud. Terrain b combines robot-sized boulders with tree debris. Terrain c includes leaves, robot-sized boulders, and weeds. Terrain d features a mix of robot-sized boulders, rocks, grass, and pine straw. Terrain e contains a dense combination of highly entangled weeds and robot-sized boulders. In these highly unpredictable environments, the feedback controller increased the robot’s speed by 30-60%, reaching up to 60% of its maximum velocity in open areas. Additional details and visual demonstrations of the outdoor experiments can be accessed in the Supplementary Information (SI) videos.
@@ VII Conclusion
VIII Discussion and Limitations
VIII-A Model limitations
In this work, we developed two probabilistic models: one capturing the correlation between the robot’s speed and contact ratio, and another predicting the contact ratio based on terrain rugosity and vertical amplitude. However, there are several limitations to these models that are worth noting and addressing in future work.
@@ VIII-A Model limitations
Our analysis in Section IV for the first model contained errors due to two key assumptions, along with other factors. First, we assumed that when a leg loses contact, the support force is lost rather than redistributed to other legs, based on the hypothesis that the force transfers to the robot’s belly. Second, we assumed the probability distribution of the slipping angle on flat terrain remains constant, though it may change on rough terrain. Additionally, sensor noise introduces errors in measuring the contact ratio, $\gamma$. To simplify calculations, we neglected the impact of collisions between the robot and the environment, despite their effect on thrust generation. These factors contributed to approximately 15% of the data points falling outside our predictions (Fig.9.B).
Further, We also assume that two models can be applied to robots with different numbers of legs. However, results in Fig.10A indicate that robots with fewer legs are more sensitive to terrain rugosity. This is primarily because our model assumes negligible yaw motion—an assumption that becomes less valid as the number of legs decreases. Observations show that when the head segment collides with an obstacle, yaw motion is induced, particularly in robots with fewer segments. For example, in the four-segment (eight-legged) configuration, yaw motion becomes significant, leading to lateral deviations and a more rapid reduction in forward speed as terrain rugosity increases from 0.17 to 0.32 (Fig.10).
To improve predictive accuracy across different robot morphologies, future work could introduce an additional term to account for leg number, modifying the speed function from $\bar{v}=f(\gamma)$ to $\bar{v}=f(\gamma)+g(n_{l})$, where $n_{l}$ represents the number of legs. This refinement would enhance the model’s applicability to robots with different configurations and improve its ability to predict locomotion performance across varied terrains.
In addition, two models in the paper rely on the contact ratio $\gamma$, without explicitly accounting for the quality of contact. In particular, the model does not differentiate between stable (firm ground) and unstable (e.g., slippery or loose) contacts, which can influence thrust generation.
For terrains with low friction coefficients, the gait framework used in this paper is designed to accommodate slipping during locomotion. In our previous work [3], we showed that thrust is governed by rate-independent Coulomb friction. During locomotion, the foot periodically slips forward and backward, balancing thrust and drag. Experimental results demonstrated that the robot’s speed (displacement per cycle) was relatively insensitive to terrain friction variations, provided there are no significant local shifts in friction. In this paper, all laboratory based experiments were conducted on surfaces with constant friction conditions. Nonetheless, we acknowledge that rapid changes in friction—such as alternating patches of mud, ice, or sand—may challenge this assumption [42], and we plan to explore these conditions in future work.
Importantly, even with friction variability, our feedback controller is expected to enhance performance by maintaining a higher contact ratio $\gamma$, resulting in greater thrust and forward speed compared to open-loop control. This is supported by our outdoor experiments (Fig. 15), where terrains (a, b, and e) included both mud and slippery boulders, and the feedback controller still achieved up to a 50% improvement in speed.
For loose granular media such as sand, where the leg-ground interaction exhibits both solid- and fluid-like behavior, we recognize that a new modeling framework will be necessary, as Coulomb-based assumptions may no longer hold. Exploring this direction is part of our ongoing and future work.
VIII-B Control framework limitations
In this paper, we developed a linear controller to improve the robot’s speed on rugged terrain; however, this simple approach has certain limitations, and there remains room for further performance enhancement.
To support our hypothesis that vertical body undulation enhances robot performance, we assume that the heights of rough terrain follow a normal distribution — a condition that aligns with our lab-based experimental setup. However, real-world terrain is typically more diverse and complex. While we demonstrated the effectiveness of our controller in more challenging outdoor tests, future work will focus on evaluating the model across a broader range of terrain types to better capture real-world variability.
@@ VIII-B Control framework limitations
We employed a sinusoidal traveling wave model to drive the robot’s vertical body undulation. While effective in mitigating environmental disturbances through amplitude adjustments, this approach cannot generate more complex vertical body shapes. Future research will investigate the impact of more discrete vertical body shape on robot performance, focusing on how subtle variations can optimize locomotion efficiency and adaptability on complex terrains. Our recent work has produced promising results [43].
@@
VIII-C Applicability to general platforms
Although the control framework is demonstrated to be effective in this paper, it remains an open question whether it is applicable to legged robots with different morphologies and dimensions.
Following this study, our group has developed additional multi-legged robots with varying numbers of legs [43], body dimensions, and leg-to-body length ratios. Across these platforms, we consistently observed that vertical body undulation enhances locomotion speed, suggesting that the benefits of this mechanism are not limited to a specific morphology. For quadrupedal and hexapod robots, we posit that introducing vertical body undulation could similarly improve performance—particularly by increasing stability on uneven or bumpy terrain. While the exact parameters may need to be tuned based on each platform’s kinematics, the underlying principle of using body undulation to mitigate environmental disturbances is expected to be broadly applicable.
Motion tracking for experiments
@@ VIII-C Applicability to general platforms
Additional experiments
-D Complaint leg test
In Section II-A, we introduced the compliant leg design, inspired by the design in [35], to increase robot’s obstacle negotiation capability. We also conducted experiments to compare the performance of rigid leg and complaint leg on rough terrain ($R_{g}=0.32$). We varied the vertical amplitude from 0 to $40^{\circ}$, with $10^{\circ}$ increment. For each condition, we performed 5 trials, and in each trial, the robot completed 3 motion cycles. The plot below presents the results of these experiments. As Fig. 16 shows, the complaint leg increases the forward speed of the robot by around 200%.
Figure 16: Robot performance with and without compliant legs. The robot completed three motion cycles on rough terrain ($R_{g}=0.32$) using both compliant and rigid leg configurations. Across all tested vertical amplitudes, the robot equipped with compliant legs achieved more than a 200% increase in speed compared to the rigid-legged configuration.
-E Gait stability at higher speeds
In the lab-based and outdoor experiments presented in this paper, the robot was operated at a temporal frequency of 1/6 Hz (6 seconds per motion cycle). To further investigate the effect of higher speeds, we conducted additional experiments at increased frequencies: 1/3 Hz (3 seconds per cycle) and 1 Hz (1 second per cycle), on rough terrain with $R_{g}=0.17$. We tested three different vertical amplitudes: $10^{\circ}$, $20^{\circ}$, and $30^{\circ}$. For each condition, we conducted five trials, with the robot completing three motion cycles per trial.
The results (Fig. 17) show that gait performance remains consistent when the temporal frequency increases from 1/6 Hz to 1/3 Hz. However, at 1 Hz, the robot’s displacement per cycle decreases to approximately 50% of the 1/6 Hz value for the smallest vertical amplitude ($10^{\circ}$), and to about 85% for the highest vertical amplitude ($30^{\circ}$). These findings suggest that larger vertical amplitudes help maintain gait performance at higher speeds. We plan to further investigate this phenomenon in future work.
Figure 17: Gait performance at higher temporal frequencies. The plot shows the average forward displacement per cycle $\bar{v}$ as a function of vertical body wave amplitude $A_{v}$, evaluated at three temporal frequencies: 1/6 Hz (6 seconds per cycle), 1/3 Hz (3 seconds per cycle), and 1 Hz (1 second per cycle), on rough terrain with $R_{g}=0.17$. Each data point represents the average over five trials, with error bars indicating standard deviation. Results show that gait performance remains stable when increasing frequency from 1/6 Hz to 1/3 Hz. At 1 Hz, performance drops significantly at low amplitudes but remains relatively high when larger vertical amplitudes are used, suggesting that greater body undulation helps maintain stability and efficiency at higher speeds.
-F $K_{p}$ sensitivity analysis
To evaluate the impact of the proportional gain parameter $k_{p}$ in our feedback controller, we conducted a series of sensitivity experiments. Recall that the vertical body wave amplitude is computed as,
$A_{v}=k_{p}(\gamma_{0}-\gamma_{s}),$
where $\gamma_{0}$ is the nominal contact ratio and $\gamma_{s}$ is the sensed contact ratio.
We varied $k_{p}$ from $0^{\circ}$ to $180^{\circ}$ in $30^{\circ}$ increments. For each value, the robot completed five motion cycles per trial, and each condition was tested across five repeated trials to ensure robustness.
When $k_{p}$ is small, the resulting vertical amplitude $A_{v}$ remains near zero, and the robot’s behavior closely resembles a gait without vertical undulation. Conversely, at higher $k_{p}$ values (approaching $180^{\circ}$), even a small difference between $\gamma_{0}$ and $\gamma_{s}$ produces a large $A_{v}$. Since we cap the vertical amplitude at $50^{\circ}$ to prevent self-collision, setting $k_{p}$ near $180^{\circ}$ effectively causes the robot to operate at its maximum vertical amplitude throughout the motion cycle.
The figure below illustrates how the average forward speed per cycle varies with $k_{p}$. As shown, speed increases with $k_{p}$, reaching a peak around $k_{p}=90^{\circ}$, after which it slightly declines. These results underscore the importance of gain selection and demonstrate the sensitivity of the controller to the $k_{p}$ parameter.
Figure 18: Sensitivity analysis of $K_{p}$. The plot shows the average forward speed per cycle $\bar{v}$ as a function of $K_{p}$, which determines the vertical body undulation amplitude according to $A_{v}=K_{p}(\gamma_{0}-\gamma_{s})$. Each data point represents the mean speed over five trials, with each trial consisting of three motion cycles. Error bars indicate the standard deviation across trials. The results show that performance improves with increasing $K_{p}$, peaking around $K_{p}=90^{\circ}$, after which the speed slightly declines. This trend highlights the importance of appropriate gain tuning in the feedback controller.
@@ Acknowledgment
We would like to thank Daniel Irvine and Jeff Aguilar for helpful discussions. We are grateful for funding from NSF STTR AWD-2335553, Army Research Office Grant AWD-001950, Georgia Research Alliance (GRA) AWD-005494, GA AIM AWD-004173 and the Dunn Family Professorship.